OpenStudy (anonymous):

Help I will fan and medal you!! The cost of a ticket to an amusement park was increased by 4 dollars. The new cost is twenty-eight dollars. a. Write a one-step equation that could be solved to find the old price. Let p represent the old price. b. Solve the equation you wrote in Part a. Show all your work. c. Explain how you know that the solution you found in Part b is the correct answer no drawlings plz

OpenStudy (texaschic101):

(a) p + 4 = 28 (b) p + 4 = 28 -- subtract 4 from both sides p = 28 - 4 p = 24 (c) check your answer by subbing in 24 for p p + 4 = 28 24 + 4 = 28 28 = 28 (correct)